Queenie Cocoa manufactures nutritious cocoa powder packaged for local market. The product is sold in boxes at RM40 per unit. The cost incurred to manufacture and market the product follows: Cost Schedule:1. Variable Costs per Box (RM) *Direct Unit=12, Direct Labour per unit=6, Manufacturing overhead=5, selling and administrative=3. 2.Fixed Costs(RM): Manufacturing overhead=300,000 and selling and administrative=120,000    

In the first year of its operation, Queenie Cocoa manufactures 50,000 boxes of which 42,000 boxes were sold. (a)   Prepare a contribution margin income statement. (b)  Calculate Queenie’s break-even point in units and ringgit. (c)   Calculate the units that Queenie must sell to achieve an after-tax profit target of RM210,000? Assume the tax rate is 40%. Indicate the margin of safety in units, at this point. (d)  Queenie wishes to incur and additional selling expense of RM35,000, other costs remain constant. Calculate the new break-even point in RM.


1
Expert's answer
2021-12-13T13:53:37-0500

a)

Contribution Margin Income Statement:


Sales = "42000\cdot40=1680000"

Variable production expenses = "50000\cdot(12+6+5+3)=1300000"

Contribution margin = Sales - Variable production expenses =

"=1680000-1300000=380000"


Fixed production expenses = "300000+120000=420000"


Net profit = Contribution margin - Fixed production expenses =

"=380000-420000=-40000"


b)

break-even point: profit = 0

x is number of uits

"40x-1300000-420000=0"

"x=1720000\/40=43000" units sold


c)

"40x(1-0.4)-1300000-420000=210000"

"x=1930000/24=80416.67\approx80417" units sold


margin of safety = (Current Sales Level – Breakeven Point) / Current Sales Level


Current Sales Level = 42000 units

for Breakeven Point:

"40x(1-0.4)-1300000-420000=0"

"x=1720000/24=71666.67\approx71667"


margin of safety = "(42000-71667)/42000=-0.71=-71\%"


d)

"40x(1-0.4)-1300000-420000-35000=0"

"x=1755000\/24=73125" units